OpenStudy (anonymous):

Why had the North and the South grown so far apart by 1861? A. The North opposed the idea that a state could leave the Union, but the South supported it. B. The South wanted to start a colony for freed slaves in Africa, but the North rejected the plan. C. The South demanded more railroads, but the North refused to pay the costs of construction. D. The North wanted to end all immigration to keep wages high, but the South needed more factory workers.

OpenStudy (anonymous):

@Tpunt1125

OpenStudy (anonymous):

This is a riddle that will answer your question @JellyCakes

OpenStudy (anonymous):

What is 5-3

OpenStudy (anonymous):

The difference + 2

OpenStudy (anonymous):

The sum minus 3

OpenStudy (anonymous):

There is your answer

OpenStudy (anonymous):

\[5-3=2\]

OpenStudy (anonymous):

\[2+2=4\]

OpenStudy (anonymous):

\[4-3=1\]

OpenStudy (anonymous):

So what is the answer 1=A 2=B 3=C 4=D

OpenStudy (anonymous):

So correct A. The North opposed the idea that a state could leave the Union, but the South supported it.